# EchoScan: Scanning Complex Indoor Geometries via Acoustic Echoes

### v1

@@ Abstract
**footnotetext: These authors contributed equally to this work.
@@ I Introduction
Digital twins have expedited innovative industrial applications spanning diverse sectors, such as navigating unfamiliar terrains or planning efficient evacuation blueprints [1]. The estimation of the indoor geometries is a crucial step in constructing realistic digital twins of indoor spaces. Traditionally, the methods for identifying indoor geometries have depended on vision-based techniques. However, achieving accurate digital representations is challenging, particularly when visual data are limited or inaccessible. Cases with insufficient visual cues, such as disaster sites or power outages, highlight the inherent limitations of vision-based approaches. Even with visible light, the estimation of occluded geometries is fundamentally challenging for vision-based digital twin construction. For such situations, acoustic indoor geometry estimation has emerged as a promising alternative that demonstrates the advantages of acoustic wave propagation through complex structures.
@@
Acoustic echoes encapsulate essential information on indoor geometrical characteristics. When the sound emitted from an audio device interacts with room boundaries, interactions such as specular and diffuse reflection, diffraction, and scattering are captured as room impulse responses (RIRs). The representative features of RIRs for indoor geometry estimation are the time-of-arrival (TOA), which represents the duration required for sound to travel from a source to a receiver, and the direction-of-arrival (DOA), which indicates the impinging direction of the reflected waves.
@@
Several data-driven methods have been proposed to overcome the limitations of conventional model-based methods that rely on low-order reflections. Deep neural networks (DNN) have been used to analyze the complex relationship between low- and high-order reflections [12, 13, 14, 15, 16, 17, 18]. Yu and Kleijn [12] used a convolutional neural network (CNN) to analyze the relationship between the RIR and room acoustic parameters and estimated the size and absorption coefficient of shoebox rooms. Poschadel et al. [13] employed a convolutional recurrent neural network (CRNN) to determine the lengths, widths, and heights of shoebox rooms using simulated RIRs. Tuna et al. [15] also utilized a CRNN architecture to infer the 2D Cartesian coordinates of a real microphone and four imaginary microphones formed by sidewalls. This data-driven method exhibited RGI performance comparable to that of the model-based technique [19] when tested with unseen measured RIRs. Despite their outstanding RGI performance, these networks can only handle shoebox rooms because of their immutable output parameter lengths. Therefore, in our previous study [16], we attempted to estimate various room geometries while disregarding the number of walls. This was possible by implementing an additional subnetwork that determines the confidence of the estimated wall parameters. However, the model cannot handle arbitrary geometries with curved walls because it estimates the coefficients of the plane-wall equation.
To circumvent these issues, this study approaches the indoor geometry estimation problem as a pixel segmentation task to infer 2D floorplan and 1D height sections of arbitrarily shaped rooms. The proposed EchoScan delivers three key contributions:

### v3

@@ I Introduction
Dgitial twins have expedited innovative industrial applications spanning diverse sectors, such as navigating unfamiliar terrains or planning efficient evacuation blueprints [1]. The estimation of the indoor geometries is a crucial step in constructing realistic digital twins of indoor spaces. The methods for identifying indoor geometries have been studied using vision-based techniques. However, achieving accurate digital representations is challenging, particularly when visual data are limited or inaccessible. Cases with insufficient visual cues, such as disaster sites or power outages, highlight the inherent limitations of vision-based approaches. Even with visible light, the estimation of occluded geometries is fundamentally challenging for vision-based digital twin construction. For such situations, acoustic indoor geometry estimation has emerged as a promising alternative that demonstrates the advantages of acoustic wave propagation through complex structures.
@@
Acoustic echoes encapsulate essential information on indoor geometrical characteristics [2]. When the sound emitted from an audio device interacts with room boundaries, interactions such as specular and diffuse reflection, diffraction, and scattering are captured as room impulse responses (RIRs). The representative features of RIRs for indoor geometry estimation are the time-of-arrival (TOA), which represents the duration required for sound to travel from a source to a receiver, and the direction-of-arrival (DOA), which indicates the impinging direction of the reflected waves.
@@
Several data-driven methods have been proposed to overcome the limitations of conventional model-based methods that rely on low-order reflections. Deep neural networks (DNN) have been used to analyze the complex relationship between low- and high-order reflections [13, 14, 15, 16, 17, 18, 19, 20]. Yu and Kleijn [13] used a convolutional neural network (CNN) to analyze the relationship between the RIR and room acoustic parameters and estimated the size and absorption coefficient of shoebox rooms. Poschadel et al. [14] employed a convolutional recurrent neural network (CRNN) to determine the lengths, widths, and heights of shoebox rooms using simulated RIRs. Tuna et al. [16] also utilized a CRNN architecture to infer the 2D Cartesian coordinates of a real microphone and four imaginary microphones formed by sidewalls. This data-driven method exhibited RGI performance comparable to that of the model-based technique [21] when tested with unseen measured RIRs. Despite their outstanding RGI performance, these networks can only handle shoebox rooms because of their immutable output parameter lengths. Therefore, in our previous studies [17, 18], we attempted to estimate various room geometries while disregarding the number of walls. This was possible by implementing an additional subnetwork that determines the confidence of the estimated wall parameters. However, the model cannot handle arbitrary geometries with curved walls because it estimates the coefficients of the plane-wall equation. To circumvent these issues, we approach the indoor geometry estimation problem as a pixel segmentation task to infer 2D floorplan and 1D height sections of arbitrarily shaped rooms.
The proposed EchoScan delivers three key contributions: